### v3

# Informativity Conditions for Multiple Signals: Properties, Experimental Design, and Applications (extended version)

Abstract
Recent studies highlight the importance of persistently exciting condition in single signal sequence for model identification and data-driven control techniques. However, maintaining prolonged excitation in control signals introduces significant challenges, as continuous excitation can reduce the lifetime of mechanical devices. In this paper, we introduce three informativity conditions for various types of multi-signal data, each augmented by weight factors. We explore the interrelations between these conditions and their rank properties in linear time-invariant systems. All three conditions can extend Willems’ fundamental lemma and are utilized to assess the properties of the system. Furthermore, we introduce open-loop experimental design methods tailored to each of the three conditions, which can synthesize the required excitation conditions either offline or online, even in the presence of limited information within each signal segment. Illustrative examples confirm that these conditions yield satisfactory outcomes in both least-squares identification and the construction of data-driven controllers.
keywords:
Informativity conditions, multiple signals, data-driven methods, experiment design
thanks: This work was supported by the National Natural Science Foundation of China under Grant 62103203. (Corresponding author: Fuyong Wang.)
1 Introduction
With the continuous development and application of data-driven technologies, the academic interest in data informativity has surged notably. In this context, the concept of persistently exciting (PE) condition (Willems et al., 2005) plays a crucial role. The PE condition requires that a signal be sufficiently rich over time to capture all dynamic characteristics of the system. This requirement is fundamental to various system identification methods (Ioannou and Fidan, 2006; Katayama et al., 2005; Markovsky et al., 2005; Markovsky, 2012; Lovera et al., 2000) and data-driven control methods (Berberich et al., 2021; Bongard et al., 2023; Liu et al., 2023; De Persis and Tesi, 2020; Park and Ikeda, 2009; Nortmann and Mylvaganam, 2023).
While persistency of excitation can indeed be enforced in theory by well-known inputs (Gevers et al., 2009) such as pseudo-random binary sequences (PRBS), broadband white noise, or periodic signals with many harmonics, their practical deployment is often constrained. First, guaranteeing PE with these signals typically requires a long single trajectory, which may not be available in experimental environments where data collection must be interrupted or restarted. Second, long-duration stochastic or high-frequency excitations can violate safety or performance constraints in many systems, e.g., process reactors or power electronic converters where sustained probing may be unsafe (Yang et al., 2024), or robotic platforms where prolonged excitation induces mechanical wear (Hamaya et al., 2021). Third, when multiple shorter trajectories are collected, naive juxtaposition of PRBS-type signals does not necessarily yield a full-rank Hankel matrix due to phase misalignment or heterogeneous signal magnitudes. These limitations highlight the need for conditions that certify data informativity beyond the classical single-trajectory PE framework.
Existing research has shown that PE condition is not a necessary prerequisite for system identification and data-driven control design (van Waarde et al., 2020b; Kang and You, 2023; Van Waarde et al., 2023). In the domain of system identification, methods such as dynamic regressor extension and mixing (Aranovskiy et al., 2016; Ortega et al., 2021) have relaxed the PE condition by extending the regression model in the time domain. Certain adaptive methods (Dhar et al., 2022) only require that excitation conditions be met during the initial phase. Furthermore, the challenge of insufficient information from a single sensor has been addressed through distributed frameworks (Chen et al., 2014; Xie and Guo, 2018) and centralized frameworks (Markovsky and Pintelon, 2015). Despite these advancements, these methods often relax the information content of system regression vectors or system trajectories, yet rarely focus on designing from the perspective of control signals. Therefore, a clear experimental design strategy for system input signals remains essential. While current experimental design theories are primarily based on the premise of satisfying identifiability (De Persis and Tesi, 2021; van Waarde, 2022; Bombois et al., 2006; Hjalmarsson, 2005), there remains a significant gap in research regarding how to design experiments that guarantee the effectiveness of identification methods when the informativeness of input signals is lacking. In such cases, the use of multiple trajectories may be necessary.
In the design of data-driven control methods, A new control framework is developed in (van Waarde et al., 2020b) that eliminates the stringent requirement of PE, but the trajectory data used still needs to guarantee the identifiability condition. The collectively persistently exciting (CPE) condition proposed in (van Waarde et al., 2020a) and (Yu et al., 2021) offers potential insights for the design of data-driven control methods using trajectories generated by insufficiently informative controllers. By organizing the Hankel matrices that evaluate the excitability of each signal sequence side-by-side into a larger rank judgment matrix, this condition extends the fundamental lemma from a single signal sequence to multiple signal sequences. However, this large matrix structure can lead to a dramatic increase in computational burden as the number of columns increases, and excessive differences in the magnitude of the values between each Hankel matrix can also cause deviations from the expected results. Additionally, it is currently an open question how to provide more choices of matrix structures for different data-driven problems (Markovsky and Dörfler, 2021).
Motivated by the aforementioned challenges, the objective of this paper is to relax the PE condition typically required for a single control sequence in data-driven methods, while offering more flexible matrix structure choices for a variety of data-driven problems. Furthermore, we aim to design experimental methodologies that ensure the effectiveness of data-driven techniques even in scenarios where data informativeness is lacking. Specifically, the contributions of this paper are summarized as follows:
  • We introduce three weighted CPE conditions for multiple signal sequences with different dimensions, same dimensions, and partially the same dimensions. All three conditions relax the PE condition requirement for a single signal sequence. Furthermore, we analyze the feasibility of implementing these conditions and their transformation relationships, and provide an initial exploration of which CPE conditions are most suitable for different problem types.
  • For the three CPE conditions, we design experimental schemes and demonstrate that these experimental designs remain effective even when signal sequence information is insufficient. We also explore how control sequences satisfying the CPE conditions guide the convergence precision of least-squares (LS) estimators. Furthermore, by utilizing the three CPE conditions, we extend Willems’ fundamental lemma, and the extended lemma yields satisfactory results in data-driven control methods.
The remainder of this paper is organized as follows. Section 2 introduces the preliminaries, covering key definitions and a review of prior work on the PE condition. In Section 3, we define three CPE conditions tied to data dimensions, analyze their transformational relationships and rank properties, and extend Willems’ fundamental lemma. Section 4 introduces experimental methods tailored to the three CPE conditions and demonstrates their validity. Section 5 introduces an identifier for single system, extending the results to the design of distributed adaptive identifiers with exponential convergence rates under CPE conditions. Section 6 provides illustrative examples of applications to validate the methodologies and theories discussed, while Section 7 concludes the paper.
2 Preliminaries
2.1 Notation
Throughout this paper, $\mathbb{Z}$, $\mathbb{N}$, $\mathbb{R}$ and $\mathbb{R}^{m\times n}$ denote the sets of integers, natural numbers, real numbers and $m\times n$ real matrices, respectively. $0_{m\times n}$ denotes a zero matrix with $m$ rows and $n$ columns. $\bm{1}_{n}$ represents the $n$-dimensional column vector with all entries set to one. $I_{n}$ and $0_{n}$ refer to the identity matrix and zero matrix of order $n$, respectively, without subscripts to indicate the suitable dimension. The modulo operation $\text{mod}(a,b)$ returns the remainder when $a$ is divided by $b$.
For a real matrix $P\in\mathbb{R}^{m\times n}$, let $P_{vec}=[P_{1},P_{2},\ldots,P_{m}]^{T}$, where $P_{i}$ represents the $i$-th row of $P$. The right kernel and left kernel of $P$ refer to the spaces of all real column vectors $v_{1}$ and row vectors $v_{2}$, respectively, satisfying $Pv_{1}=0$ and $v_{2}P=0$. The image, right kernel, left kernel, rank and Moore–Penrose pseudoinverse of $P$ are denoted as $\text{im}\,P$, $\text{ker}\,P$, $\text{leftker}\,P$, $\text{rank}\,P$ and $P^{{\dagger}}$, respectively. Additionally, let $\lambda_{max}(P)$, $P||$, and $P||_{2}$ represent the maximal eigenvalue, Frobenius norm, and spectral norm of $P$, respectively.
Given a signal $z$: $\mathbb{Z}\rightarrow\mathbb{R}^{m}$, we define $z^{[k,k+T_{0}]}$ as
$z^{[k,k+T_{0}]}=\begin{bmatrix}{z(k)}\\ \vdots\\ {z(k+{T_{0}})}\end{bmatrix}$,
where $k\in\mathbb{Z}$ and $T_{0}\in\mathbb{N}$. The block Hankel matrix of order $L$ associated to $z^{[k,k+T_{0}]}$ is defined as
${H_{L}}({z^{[k,k+{T_{0}}]}})$
$~~=\begin{bmatrix}{z(k)}&{z(k+1)}&\cdots&{z(k+{T_{0}}-L+1)}\\ {z(k+1)}&{z(k+2)}&\cdots&{z(k+{T_{0}}-L+2)}\\ \vdots&\vdots&\ddots&\vdots\\ {z(k+L-1)}&{z(k+L)}&\cdots&{z(k+{T_{0}})}\end{bmatrix}$.
2.2 Persistency of Excitation for a Single Sequence
The concept of persistency of excitation for a single signal sequence, as defined in Willems et al. (2005), is elucidated below.
Definition 1.
(Persistently Exciting Condition (Willems et al., 2005)). A signal sequence $z^{[0,T_{0}-1]}$ with $z(k)\in\mathbb{R}^{m}$, $k=0,...,T_{0}-1$ is PE of order $L$ if the block Hankel matrix $H_{L}(z^{[0,T_{0}-1]})$ has full row rank $mL$.
It is intuitive from this definition that a PE signal of order $L$ possesses the ability to span the entire real space $\mathbb{R}^{mL}$. In particular, if the input signal to a controllable linear time-invariant (LTI) system exhibits sufficiently high-order persistency of excitation, some attractive properties arise. Consider the LTI system
$x(k+1)=Ax(k)+Bu(k),$ (1)
where $A\in{\mathbb{R}^{n\times n}}$, ${B}\in{\mathbb{R}^{n\times{m}}}$, ${x(k)}\in{\mathbb{R}^{n}}$ and ${u(k)}\in{\mathbb{R}^{m}}$ represent the state vector and control input vector, respectively. The pair $(A,B)$ is controllable. A fundamental property in (Willems et al., 2005, Cor. 2) demonstrates that the LTI system (1) satisfies the rank condition
$\text{rank}\,\begin{bmatrix}{{H_{1}}({x^{[0,{T_{0}}-L]}})}\\ {{H_{L}}({u^{[0,{T_{0}}-1]}})}\end{bmatrix}=n+mL$ (2)
if the control input $u^{[0,{T_{0}}-1]}$ is PE of order $L+n$.
In addition, the rank condition (2) subtly implies that the sequences ${\bar{x}}^{[0,L-1]}$}} and ${\bar{u}}^{[0,L-1]}$}} form an $L$-long state-input trajectory of (1) if and only if there exists $g\in\mathbb{R}^{T_{0}-L+1}$ such that
$\begin{bmatrix}{{{\bar{x}}^{[0,L-1]}}}\\ {{{\bar{u}}^{[0,L-1]}}}\end{bmatrix}=\begin{bmatrix}{{H_{L}}({x^{[0,{T_{0}}-1]}})}\\ {{H_{L}}({u^{[0,{T_{0}}-1]}})}\end{bmatrix}g.$ (3)
This property indicates that the image of the matrix $\begin{bmatrix}{{H_{L}}({x^{[0,{T_{0}}-1]}})}\\ {{H_{L}}({u^{[0,{T_{0}}-1]}})}\end{bmatrix}$ is the representation of the $L$-long state-input trajectory space of the LTI system (1). It has been rigorously proved in the behavioral system framework (Willems et al., 2005, Th. 1) and the state-space framework (De Persis and Tesi, 2020, Lem. 2), respectively.
The properties described by (2) and (3), often referred to as Willems’ fundamental lemma due to their foundational importance for system identification and data-driven control.
Indeed, if the input sequence $u^{[0,{T_{0}}-1]}$ of system (1) is PE of order $n+1$, the matrix $\begin{bmatrix}{{H_{1}}({x^{[0,{T_{0}}-1]}})}\\ {{H_{1}}({u^{[0,{T_{0}}-1]}})}\end{bmatrix}$ has full row rank. This condition enables the recovery of the system matrices $A$ and $B$ from the data $x^{[0,{T_{0}}-1]}$ and $u^{[0,{T_{0}}-1]}$, i.e., identifiability condition (see (Markovsky and Dörfler, 2023; van Waarde et al., 2020b)). The system matrices can be uniquely determined by solving the following equation:
$\begin{bmatrix}A&B\end{bmatrix}={H_{1}}({x^{[1,{T_{0}}]}}){\begin{bmatrix}{{H_{1}}({x^{[0,{T_{0}}-1]}})}\\ {{H_{1}}({u^{[0,{T_{0}}-1]}})}\end{bmatrix}^{\dagger}}$. (4)
On the other hand, under the influence of the prior control input $u^{[0,T_{0}-1]}$ that satisfies the PE condition, the system trajectories ${\bar{x}}^{[0,L-1]}$ and ${\bar{u}}^{[0,L-1]}$ can form the non-parametric representation in (3), which solely relies on the prior data. Consequently, in addition to its important role in system identification, PE condition has received widespread attention in the development of data-driven controllers such as parametric model reconstruction methods (De Persis and Tesi, 2020; Nortmann and Mylvaganam, 2023) and data-driven model predictive control methods (Berberich et al., 2021; Bongard et al., 2023; Liu et al., 2023).
Even though it has been shown how to ensure the informativeness of a single control sequence (Gevers et al., 2009), in some cases it may still not be possible to generate sufficiently informative trajectories due to the many limitations of the experiment. However, designing experiments with insufficiently informative controllers fail to ensure the identifiability condition and the feasibility of data-driven approaches.
Example 1.
When applying a feedback controller $u=Kx$, where $K$ is the control gain, we have
${{H_{L}}({u^{[0,{T_{0}}-1]}})}=\begin{bmatrix}{K{{(A+BK)}^{0}}{H_{1}}({x^{[0,{T_{0}}-L]}})}\\ \vdots\\ {K{{(A+BK)}^{L-1}}{H_{1}}({x^{[0,{T_{0}}-L]}})}\end{bmatrix}$,
which means that every row of ${H_{L}}({u^{[0,{T_{0}}-1]}})$} is in the row space of ${H_{1}}({x^{[0,{T_{0}}-L]}})$. Therefore,
$\text{rank}\ {{H_{L}}({u^{[0,{T_{0}}-1]}})}<mL.$
Consequently,
$\text{rank}\,\begin{bmatrix}{{H_{1}}({x^{[0,{T_{0}}-L]}})}\\ {{H_{L}}({u^{[0,{T_{0}}-1]}})}\end{bmatrix}\leq n$
and rank condition (2) is not satisfied. In this case, neither identifiability nor fundamental lemma can be guaranteed.
This raises three natural questions. First, can multiple non-persistently exciting sequences be employed to relax the requirement for persistent excitation of a single sequence as defined in Definition 1, and if so, how can they be utilized? Second, how can we design experiments to address the first question? Third, is it feasible to leverage multiple non-persistently exciting control sequences for model identification and data-driven controller design? These three questions are the subject of the remainder of this paper.
3 Three Collectively Persistently Exciting Conditions and Their Properties
In this section, we introduce three collectively persistently exciting (CPE) conditions for multi-trajectory data, aiming to eliminate the requirement of persistently excited single-trajectory sequence as specified in (2). These conditions are delineated from three perspectives: same-dimension data, different-dimension data, and partially same-dimension data. Theorem 1 elucidates the transformation relations among these CPE conditions, while Lemma 1 establishes a rank condition akin to (2), and Lemma 2 extends Willems’ fundamental lemma under CPE conditions.
Figure 1: Transformative relationship between the three CPE conditions. The left side of the figure illustrates the synthesis of the conditions, while the right side presents the specific conditions. The arrows denote the derived relationships.
3.1 Three Collectively Persistently Exciting Conditions
In this subsection, we delineate the CPE conditions into three types: cumulative, mosaic, and hybrid, tailored for multi-trajectory sequences. Unlike the traditional PE condition, where each trajectory needs to exhibit sufficient excitation and satisfy identifiable conditions, our approach shifts focus to the collective impact. The three distinct matrix structures afford greater flexibility for trajectories of varying lengths, thereby catering to a wide array of problem types.
Before proceeding, let $\{z_{i}^{[k,k+T_{i}]}\}_{i=1}^{p}$ represent the sequences $z_{1}^{[k,k+T_{1}]},z_{2}^{[k,k+T_{2}]},\ldots,z_{p}^{[k,k+T_{p}]}$, where $z_{i}^{[k,k+T_{i}]}$ denotes the value of the $i$-th signal $z_{i}$ within a specific time interval $k,k+T_{i}]$, with $T_{i}$ representing a natural number associated with index $i$.
We begin by introducing a mosaic collectively persistently exciting (MCPE) condition for signal sequences of different dimensions, which is shown to be the easiest exciting condition to satisfy in the subsequent discussion (see Theorem 1).
Definition 2.
(Collectively Persistently Exciting Condition (Type-I): mosaic). Multiple signals $\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p}$ with $z_{i}(k)\in\mathbb{R}^{m}$, $k=0,1,...,{T_{i}-1}$ are mosaic collectively persistently exciting of order $L$ if the matrix
${H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}$
$~~~~~~~~~~~~~~~~~~~={\begin{bmatrix}{{\alpha_{1}}{H_{L}}({z_{1}^{[0,{T_{1}}-1]}})}\ \ldots\ {{\alpha_{p}}{H_{L}}({z_{p}^{[0,{T_{p}}-1]}})}\end{bmatrix}}$
has full row rank $mL$ for any $\alpha_{i}\neq 0$.
Compared to the PE condition for a single trajectory, the MCPE condition is motivated by practical scenarios where generating one sufficiently long informative trajectory is infeasible. Typical examples include process systems constrained by safety limits (e.g., reactors or power electronics that cannot be excited for long durations), robotic platforms where prolonged excitation causes mechanical wear, or networked systems with limited storage or communication. In such cases, multiple short (non-PE) trajectories are far easier to obtain.
The MCPE condition does not specify the number of dimensions of each signal sequence, i.e., the $T_{i}$ of each signal can be different. Only the sum of all signal lengths is required to satisfy $\sum\limits_{i=1}^{p}{{T_{i}}}\geq mL+p(L-1)$. However, this flexibility results in an unrestricted number of columns in the rank judgment matrix $H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})$}, potentially leading to an excessive number of redundant columns in the matrix. This abundance of columns can significantly increase computational complexity in practical applications such as model predictive control Berberich et al. (2021); Bongard et al. (2023); Liu et al. (2023).
Therefore, in the following definition, we introduce a cumulative collectively persistently exciting (CCPE) condition for same-dimensional signal sequences. This condition is subsequently confirmed to be the most challenging to satisfy but is the most suitable in terms of computational complexity for application in the design of data-driven control methods.
Definition 3.
(Collectively Persistently Exciting Condition (Type-II): cumulative). Multiple signals $\{z^{[0,T_{0}-1]}_{i}\}_{i=1}^{p}$ with $z_{i}(k)\in\mathbb{R}^{m}$, $k=0,1,...,{T_{0}-1}$ are cumulative collectively persistently exciting of order $L$ if the matrix
${H_{L}^{cum}(\{z^{[0,T_{0}-1]}_{i}\}_{i=1}^{p})}=\sum\limits_{i=1}^{p}{{\alpha_{i}}{H_{L}}({z^{[0,T_{0}-1]}_{i}})}$
has full row rank $mL$ for any $\alpha_{i}\neq 0$.
The CCPE condition strictly demands that all signal sequences share the same dimension/length, with a minimum requirement of ${T_{0}}\geq(m+1)L-1$. This stringent dimensionality requirement for different signal sequences ensures that the rank-judgment matrix $H_{L}^{cum}(\{z^{[0,T_{0}-1]}_{i}\}_{i=1}^{p})$} can be confined to a relatively smaller size, specifically, $mL\times{(T_{0}-L+1)}$ dimensions. Consequently, when it comes to computations using rank judgment matrices, the CCPE condition can better restrict the complexity of the matrices compared to the MCPE condition. This restraint effectively serves the purpose of reducing computational costs.
Next, we introduce a hybrid collectively persistently exciting (HCPE) condition for partially same-dimensional signal sequences. This condition employs a rank judgment matrix that combines the matrix structures from both the CCPE condition and the MCPE condition, leveraging the advantages and mitigating the disadvantages of each.
Definition 4.
(Collectively Persistently Exciting Condition (Type-III): hybrid). Multiple signals $\{z^{[0,T_{0}-1]}_{i}\}_{i=1}^{\bar{p}}$ and $\{z^{[0,T_{i}-1]}_{i}\}_{i=\bar{p}+1}^{p}$ with $z_{i}(k)\in\mathbb{R}^{m}$, $k=0,1,...,{T_{i}-1}$ are hybrid collectively persistently exciting of order $L$ if the matrix
${H_{L}^{hyb}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}$
$~~~~~~={\begin{bmatrix}{H_{L}^{cum}(\{z^{[0,T_{0}-1]}_{i}\}_{i=1}^{\bar{p}})}&{H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=\bar{p}+1}^{p})}\end{bmatrix}}$
has full row rank $mL$ for any $\alpha_{i}\neq 0$.
The HCPE condition requires the first $\bar{p}$ signal sequences (which can be any $\bar{p}$ sequences, without loss of generality, and are taken here as the first $\bar{p}$) to be of the same dimension/length, with $T_{0}+\sum\limits_{i=\bar{p}}^{p}{{T_{i}}}\geq mL+(p-\bar{p}+1)(L-1)$.
It is important to note that none of the three CPE conditions require each signal sequence to be PE. Rather, they allow for insufficient information sequences but mandate that the overall behavior demonstrates a PE-like performance, which is physically achievable. In Section 4, we propose experimental methods to synthesize the three CPE conditions in cases where each signal segment is not PE.
The benefit of MCPE lies not in reducing the number of matrix columns, but in simplifying the overall computational procedure of informativity verification while enhancing flexibility in handling heterogeneous trajectories. Instead of repeatedly checking PE on each trajectory or artificially aligning all trajectories to the same horizon (as required by CCPE), MCPE organizes all heterogeneous trajectories into a single rank test. Thus, only one unified check is needed, while weighting factors mitigate numerical ill-conditioning caused by magnitude discrepancies across datasets. Although the resulting matrix may be wider, the computation pipeline is considerably simpler in certain scenarios, especially when handling heterogeneous-length data, which is shown with the illustrative example in the Section 6.5.
Remark 1.
The CPE condition proposed in van Waarde et al. (2020a) shares similarities with the first type introduced in this paper, the MCPE condition, as both combine Hankel matrices in a side-by-side fashion. However, the MCPE condition distinguishes itself by incorporating weight factors, $\alpha_{i}$, which allows it to perform better when handling data with significant order-of-magnitude differences. Additionally, a careful selection of the weighting factors can enhance both performance and computational efficiency when employing data-driven methods, as demonstrated in the examples in Section 6.2. Notably, despite the introduction of weight factors, the difficulty of synthesis remains unchanged. This is attributed to the fact that for any $\alpha_{i}\neq 0$, the rank of the matrix $H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})$} remains the same as when $\alpha_{i}=1$.
3.2 Transformation Relations and rank properties of CPE Conditions
In this subsection, we analyze the transformation relations between the three CPE conditions. The result reflects the difficulty of realizing the three conditions, and provides some theoretical basis for the selection of the three conditions in practical identification and control problems.
Theorem 1.
For multiple signals $\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p}$ with $z_{i}(k)\in\mathbb{R}^{m}$, $k=0,1,...,{T_{i}-1}$, the following transformations hold:
  • $1).$
    If $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are CCPE $\Longrightarrow$ $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also MCPE and HCPE;
  • $2).$
    If $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are MCPE and satisfy
    • i).
      $T_{0}=T_{1}=...=T_{p}$, $\text{im}\,{{H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}^{T}}\cap\text{leftker}\,{\textbf{1}_{p}}\otimes{I_{{T_{0}}-L+1}}=\{0\}$ $\Longrightarrow$ $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also CCPE;
    • ii).
      $T_{0}=T_{1}=...=T_{\bar{p}}$, $\text{im}\,{{H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{\bar{p}})}^{T}}\cap\text{leftker}\,{\begin{bmatrix}{{\textbf{1}_{\bar{p}}}}&{{0_{\bar{p}\times(p-\bar{p})}}}\\ {{0_{p-\bar{p}}}}&{{I_{p-\bar{p}}}}\end{bmatrix}}\otimes{I_{{T_{0}}-L+1}}=\{0\}$ $\Longrightarrow$ $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also HCPE;
  • $3).$
    If $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are HCPE,
    • i).
      $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also MCPE;
    • ii).
      $T_{0}=T_{\bar{p}+1}=...=T_{p}$, $\text{im}\,{{H_{L}^{hyb}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}^{T}}\cap\text{leftker}\,{\textbf{1}_{\bar{p}+1}}\otimes{I_{{T_{0}}-L+1}}=\{0\}$ $\Longrightarrow$ $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also CCPE.
Proof. See the Appendix. A. $\blacksquare$
Theorem 1 demonstrates that the MCPE condition is the most straightforward to achieve, followed by the HCPE condition, with the CCPE condition being the most difficult. A significant finding is that satisfying both the CCPE and HCPE conditions results in fulfilling the MCPE condition. Fig. 1 illustrates the transformation relationship among the three conditions.
In the following lemma, we establish the rank condition under three CPE conditions, corresponding to the rank condition (2). This lemma is pivotal in extending Willems’ fundamental lemma (see Lemma 2).
Lemma 1.
Assuming $\{u_{i}^{[0,{T_{i}-1}]},x_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ constitutes the $p$-segment input-state trajectory of the system (1).
  • $1).$
    If $T_{0}=T_{1}=T_{2}=...=T_{p}$ and $\{u_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}$ are CCPE of order $L+n$, then
    $\text{rank}\begin{bmatrix}{H_{1}^{cum}(\{{x_{i}^{[0,{T_{i}-L}]}\}_{i=1}^{p}})}\\ {H_{L}^{cum}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix}=n+mL.$
  • $2).$
    If $\{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are MCPE of order $L+n$, then
    $\text{rank}\begin{bmatrix}{H_{1}^{mos}(\{{x_{i}^{[0,{T_{i}-L}]}\}_{i=1}^{p}})}\\ {H_{L}^{mos}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix}=n+mL.$
  • $3).$
    If $T_{0}=T_{1}=T_{2}=...=T_{\bar{p}}$ and $\{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are HCPE of order $L+n$, then
    $\text{rank}\begin{bmatrix}{H_{1}^{hyb}(\{{x_{i}^{[0,{T_{i}-L}]}\}_{i=1}^{p}})}\\ {H_{L}^{hyb}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix}=n+mL.$
Proof. See the Appendix. B. $\blacksquare$
3.3 Extension of Fundamental Lemma
In this subsection, we delve into the fundamental lemmas under the CPE condition, extending the traditional Willems’ lemma.
For a single-input signal sequence satisfying the PE condition of order $L+n$, (3) always holds. A similar conclusion applies when the signal sequence extends to more than one.
Lemma 2.
Assuming $\{u_{i}^{[0,{T_{i}-1}]},x_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ constitutes the $p$-segment input-state trajectory of the system (1). Then, the following holds.
  • $1).$
    If $T_{0}=T_{1}=T_{2}=...=T_{p}$ and $\{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are CCPE of order $L+n$, then any $L$-long input-state trajectory of system (1) can be expressed as
    $\begin{bmatrix}{H_{L}^{cum}(\{{x_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\\ {H_{L}^{cum}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix}g,$
    where $g$ is a real vector.
  • $2).$
    If $\{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are MCPE of order $L+n$, then any $L$-long input-state trajectory of system (1) can be expressed as
    $\begin{bmatrix}{H_{L}^{mos}(\{{x_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\\ {H_{L}^{mos}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix}g,$
    where $g$ is a real vector.
  • $3).$
    If $T_{0}=T_{1}=T_{2}=...=T_{\bar{p}}$ and $\{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are HCPE of order $L+n$, then any $L$-long input-state trajectory of system (1) can be expressed as
    $\begin{bmatrix}{H_{L}^{hyb}(\{{x_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\\ {H_{L}^{hyb}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix}g,$
    where $g$ is a real vector.
Proof. By the dynamics of system (1), one has
$\begin{bmatrix}{H_{L}^{cum}(\{{u_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})}\\ {H_{L}^{cum}(\{{x_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})}\end{bmatrix}g$
$~~~~~~~~~~~~~~=\begin{bmatrix}I&0\\ \mathcal{T}_{L}&\mathcal{O}_{L}\end{bmatrix}\begin{bmatrix}{H_{L}^{cum}(\{{u_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})}\\ {H_{1}^{cum}(\{{x_{i}^{[0,{T_{0}-L}]}\}_{i=1}^{p}})}\end{bmatrix}g,$ (5)
where $\mathcal{T}_{L}$ and $\mathcal{O}_{L}$ are defined as
${{\mathcal{T}}_{L}}=\begin{bmatrix}0&0&0&\cdots&0\\ B&0&0&\cdots&0\\ {AB}&B&0&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ {{A^{L-2}}B}&{{A^{L-2}}B}&{{A^{L-2}}B}&\cdots&0\end{bmatrix}$
${{\mathcal{O}}_{L}}={\begin{bmatrix}I&{{A^{T}}}&{{{({A^{2}})}^{T}}}&\cdots&{{{({A^{L-1}})}^{T}}}\end{bmatrix}^{T}}.$
Since $\begin{bmatrix}{H_{L}^{cum}(\{{u_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})}\\ {H_{1}^{cum}(\{{x_{i}^{[0,{T_{0}-L}]}\}_{i=1}^{p}})}\end{bmatrix}$ has full row rank by Lemma 1, all vectors consisting of the initial state $\bar{x}(0)$ and the $L$-long input sequence $\bar{u}^{[0,L-1]}$ of system (1) can be represented by a linear combination of the columns of $\begin{bmatrix}{H_{L}^{cum}(\{{u_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})}\\ {H_{1}^{cum}(\{{x_{i}^{[0,{T_{0}-L}]}\}_{i=1}^{p}})}\end{bmatrix}$, i.e.,
$\begin{bmatrix}{H_{L}^{cum}(\{{u_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})}\\ {H_{1}^{cum}(\{{x_{i}^{[0,{T_{0}-L}]}\}_{i=1}^{p}})}\end{bmatrix}g=\begin{bmatrix}{{{\bar{u}}^{[0,L-1]}}}\\ {{{\bar{x}}(0)}}\end{bmatrix}$.
Therefore, (3.3) can be transformed into
$\begin{bmatrix}{H_{L}^{cum}(\{{u_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})}\\ {H_{L}^{cum}(\{{x_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})}\end{bmatrix}g$ $=\begin{bmatrix}I&0\\ {{{\mathcal{T}}_{L}}}&{{{\mathcal{O}}_{L}}}\end{bmatrix}\begin{bmatrix}{{{\bar{u}}^{[0,L-1]}}}\\ {{{\bar{x}}(0)}}\end{bmatrix}$
$=\begin{bmatrix}{{{\bar{u}}^{[0,L-1]}}}\\ {{{\bar{x}}^{[0,L-1]}}}\end{bmatrix}$,
which leads to conclusion 1). The proofs of conclusions 2) and 3) follow a similar process and are omitted here.
$\blacksquare$
Similar to the original fundamental lemma, Lemma 2 establishes that $L$-long trajectories can be expressed as linear combinations of columns of certain matrices. It maintains reliance on excitation conditions (CCPE, MCPE, HCPE) to ensure the validity of the representation. However, this lemma explicitly addresses $p$-segment input-state trajectories, relaxes the PE condition, and allows the excitation condition to hold collectively across all segments rather than individually for each segment.
By introducing CCPE, MCPE, and HCPE, the extended lemma formalizes the conditions under which segmented data collectively provide sufficient information for trajectory representation. This generalization broadens the lemma’s applicability to scenarios such as systems with distributed sensing, segmented experiments, or partial data. Notably, the work in van Waarde et al. (2020a) marks an initial step toward addressing multiple trajectories in Willems’ Lemma, while Lemma 2 strictly extends and incorporates the results from van Waarde et al. (2020a). In special or complex cases (e.g., significant numerical discrepancies between trajectories or equal trajectory lengths), the additional flexibility afforded by Lemma 2 enhances its applicability.
Remark 2.
Willems’ fundamental lemma has become a cornerstone in the development of data-driven control methodologies. Lemmas 1 and 2 share core properties with Williams’ lemma, positioning them as fundamental tools for similar applications. Specifically, Lemma 1 enables results analogous to (De Persis and Tesi, 2020, Th. 3), where the problem of stabilizing controller design is reformulated as a linear matrix inequality (LMI) optimization task. Additionally, Lemma 2 supports the construction of predictive models, providing a foundation for data-driven model predictive control (MPC) designs, as demonstrated in works such as Berberich et al. (2021); Bongard et al. (2023); Liu et al. (2023). Therefore, Lemmas 1 and 2 extend the applicability of data-driven techniques to scenarios involving multiple a priori trajectories, offering broader prospects for practical applications. Moreover, while the identification methods and extension lemma presented in this paper are based on fully observable LTI systems, they can be easily extended to input-output state-space systems, as the core application of these methods is rooted in Lemma 1 , which holds for input-output systems as well.
Remark 3.
Both applications discussed above rely on pre-collected data and employ non-adaptive methods. In reality, all three informativeness conditions can be reformulated to resemble the common assumptions about signals in adaptive methods. Specifically, since $H_{L}^{mos}(\{z_{i}^{[0,T_{i}-1]}\}_{i=1}^{p})$ has full row rank, it follows that
$H_{L}^{mos}(\{z_{i}^{[0,T_{i}-1]}\}_{i=1}^{p})\left(H_{L}^{mos}(\{z_{i}^{[0,T_{i}-1]}\}_{i=1}^{p})\right)^{T}>0.$
Therefore, there exists a constant $\beta_{0}>0$ such that
$\sum_{i=1}^{p}\sum_{j=0}^{L-1}\alpha_{i}^{2}\phi_{i,j}\phi_{i,j}^{T}>\beta_{0}I,$
where $\phi_{i,j}=\left[z_{i}(j)\quad z_{i}(j+1)\quad\cdots\quad z_{i}(j+T_{i}-L)\right].$
This condition is analogous to the informativeness condition in adaptive methods. According to Theorem 1, the same holds for both the CCPE and HCPE conditions. Consequently, the informativeness condition defined in this paper has significant potential for application to adaptive data-driven methods as well. Detailed results regarding the three conditions in distributed adaptive system identification methods are elaborated in the Section 5.
4 Experimental Design for CPE Conditions
In this section, we propose the experimental design approach to fulfill the three CPE conditions. This methodology is motivated by two critical challenges in practical applications. 1) Overcoming the limitation of single-trajectory PE requirements: Traditional PE conditions demand that individual trajectories exhibit sufficient richness (i.e., full rank conditions), which can be restrictive or infeasible in real-world scenarios (e.g., sparse or intermittent data). By contrast, our approach relaxes this stringent requirement, allowing non-PE signals to collectively achieve CPE through strategic design. 2) Unified online/offline applicability: Many existing methods are limited to either offline (pre-recorded data) or online (real-time) settings, but not both. Our framework is deliberately designed for dual-mode compatibility, ensuring flexibility in deployment.
Signal design for the MCPE condition: For the MCPE condition, the design ensures that the matrix $H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})$} takes the following form:
Each group of $L$ columns forms an approximately upper triangular structure, with the diagonal elements of the first $m$ submatrices being linearly independent. To construct the desired matrix using $p$ signals, it is necessary to identify which of the first $L$ values of each signal appears as diagonal elements in the submatrices, i.e., $z_{i}(\Delta(i))$, where $\Delta_{1}=L-1$, $\Delta_{i}=L-1-\text{mod}(\sum^{i-1}_{j=1}(T_{j}-L+1),L)$ for $i>1$.
If $T_{i-1}\geq\Delta_{i-1}+L$, then a submatrix spans from $z_{i-1}(\Delta(i-1))$ to $z_{i}(\Delta(i))$, and $z_{i}(\Delta(i))$ must be designed such that $z_{i}(\Delta_{i})\notin\text{im}\ [z_{1}(\Delta_{1})\cdots z_{i-1}(\Delta_{i-1})]$. Otherwise, the submatrix is not spanned, and $z_{i}(\Delta(i))$ is chosen as $z_{i}(\Delta_{i})\in\text{im}\ z_{i-1}(\Delta_{i-1})$.
For k>L1k>L-1, if $z_{i}(k)$ corresponds to the start of the diagonal of a submatrix, i.e., $T_{i}^{s}+k\in\{2L-1,3L-1,\ldots,(m+1)L-1\}$, where $T_{1}^{s}=0$, and $T_{i}^{s}=\sum^{i-1}_{j=1}(T_{j}-L+1)$ for $i>1$, it is necessary to verify whether $z_{i}$ can still form $L$ columns. If feasible, $z_{i}(k)$ is designed to remain linearly independent from the existing diagonal elements. If this is not feasible, then $z_{i}(k)$ need to satify $z_{i}(k)\in\text{im}\ z_{i+1}(\Delta_{i+1})$. The detailed design process is shown below:
Consider multiple signals $\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p}$, where $z_{i}(k)\in\mathbb{R}^{m}$ for $k=0,1,...,{T_{i}-1}$. Without loss of generality, assume that $L\leq T_{1}\leq T_{2}\leq...\leq T_{p}<(m+1)L-1$.
Construction of $z_{i}(0),z_{1}(1),\ldots,z_{i}(L-1)$ for $i=1,\ldots,p$:
  • For $i=1$:
    Select $z_{i}^{[0,L-2]}$ arbitrarily, and set $\text{rank}\ z_{i}(L-1)=1$.
  • For $i>1$:
    Select $z_{i}^{[0,\Delta_{i}-1]}$ arbitrarily for $\Delta_{i}>1$, and choose $z_{i}(\Delta_{i})$ as follows:
[…]
    Then, set $z^{[\Delta_{i}+1,L-1]}_{i}=0$ for $\Delta_{i}<L-1$.
Construction of $z_{i}(L),z_{i}(L+1),\ldots,z_{i}(T_{i}-1)$ for $i=1,\ldots,p$:
  • For $k<T_{i}$ and $T_{i}^{s}+k\in\{2L-1,3L-1,\ldots,(m+1)L-1\}$:
    Select $z_{i}(k)$ as follows:
    $\begin{cases}z_{i}(k)\notin\text{im}\ [z_{1}\left(\Delta_{1}\right)\cdots z_{p}(\Delta_{p})\ \bar{z}_{1}\cdots\bar{z}_{p}],&\text{if}\ T_{i}-k\geqslant L,\\ z_{i}\left(k\right)\in\text{im}\ z_{i+1}\left(\Delta_{i+1}\right),&\text{otherwise},\end{cases}$ where $\bar{z}_{i}=[z_{i}(L-1)\ z_{i}(2L-1)\cdots]$.
  • For $k<T_{i}$ and $T_{i}^{s}+k\notin\{2L-1,3L-1,\ldots,(m+1)L-1\}$:
    Set $z_{i}(k)=0$.
Signal design for the CCPE condition: The idea behind the design of the CCPE condition is also to construct the first $mL$ columns of $H_{L}^{cum}(\{z^{[0,T_{0}-1]}_{i}\}_{i=1}^{p})$} to be a matrix similar to the previous one. The detailed design process is shown below:
Consider multiple signals $\{z^{[0,T_{0}-1]}_{i}\}_{i=1}^{p}$, where $z_{i}(k)\in\mathbb{R}^{m}$ for $k=0,1,...,{T_{0}-1}$. Define $Z_{i}^{j}=\left[z_{i}\begin{matrix}\left(j\right)&z_{i}\left(j+L\right)&\cdots&z_{i}(j+\left(m-1)L\right)\end{matrix}\right]$.
Construction of $z_{i}(0),...,z_{i}(mL-1)$ for $i=1,\ldots,p$:
  • Choose $z_{i}(k)$ such that $\text{rank}\ \sum^{p}_{i=1}\alpha_{i}z_{i}\left(k\right)=1$ for $k\in\{L-1,2L-1,\ldots,mL-1\}$, and $\sum^{p}_{i=1}\alpha_{i}z_{i}\left(k\right)=0$ otherwise.
  • Ensure that
    $\sum^{p}_{i=1}\alpha_{i}\left[z_{i}\begin{matrix}\left(L-1\right)&z_{i}\left(2L-1\right)&\cdots&z_{i}\left(mL-1\right)\end{matrix}\right]$
    has full rank.
  • Ensure $\text{rank}\ Z_{i}^{0}=m$ for $i=1,\ldots,p$.
Construction of $z_{i}(k),k>mL-1$ for $i=1,\ldots,p$:
  • Design $z_{i}\left(k\right)$ as
    $z_{i}\left(k\right)=Z^{(k-mL+1)}_{i}({Z^{0}_{i})}^{-1}z_{i}\left(mL-1\right).$ (6)
Signal design for the HCPE condition: The design idea for the HCPE condition combines the MCPE and CCPE conditions as shown below:
Consider multiple signals $\{z^{[0,T_{0}-1]}_{i}\}_{i=1}^{\bar{p}}$ and $\{z^{[0,T_{i}-1]}_{i}\}_{i=\bar{p}+1}^{p}$, where $z_{i}(k)\in\mathbb{R}^{m}$, $k=0,1,\ldots,{T_{i}-1}$. Without loss of generality, assume that $L\leq T_{\bar{p}+1}\leq T_{\bar{p}+2}\leq...\leq T_{p}<(m+1)L$.
Construction of $z_{i}(0),\ldots,z_{i}(T_{0}-1)$ for $i=1,\ldots,\bar{p}:$
  • If $T_{0}\geqslant(m+1)L$:
    Use the CCPE signal design method.
  • If $T_{0}<(m+1)L$:
    Choose $z_{i}(k)$ such that $\text{rank}\ \sum^{\bar{p}}_{i=1}\alpha_{i}z_{i}\left(k\right)=1$ for $k\in\{L-1,2L-1,\ldots,mL-1\}$, and $\sum^{\bar{p}}_{i=1}\alpha_{i}z_{i}\left(k\right)=0$ otherwise. Ensure that
    $\sum^{\bar{p}}_{i=1}\alpha_{i}\left[z_{i}\begin{matrix}\left(L-1\right)&z_{i}\left(2L-1\right)&\cdots\end{matrix}\right]$
    has full column rank.
Construction of $z_{i}(0),\ldots,z_{i}(T_{i}-1)$ for $i=\bar{p}+1,\ldots,p$:
  • Let $\sum^{\bar{p}}_{i=1}\alpha_{i}z_{i}$ represent the initial signal, with $\{z_{i}\}_{i=\bar{p}+1}^{p}$ serving as the subsequent signals. Employ the MCPE signal design method to construct $z_{i}(0),\ldots,z_{i}(T_{i}-1)$ for $i=\bar{p}+1,\ldots,p$.
Theorem 2.
For multiple signals $\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p}$, where $z_{i}(k)\in\mathbb{R}^{m}$ for $k=0,1,\ldots,{T_{i}-1}$.
  • $1).$
    Using the MCPE design method, the signal length required to satisfy $\text{rank}\ {H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}=mL$ is $\sum\limits_{i=1}^{p}{{T_{i}}}\geq mL+p(L-1)$.
  • $2).$
    Using the CCPE design method, the signal length required to satisfy $\text{rank}\ {H_{L}^{cum}(\{z^{[0,T_{0}-1]}_{i}\}_{i=1}^{p})}=mL$ is ${T_{0}}\geq(m+1)L-1$.
  • $3).$
    Using the HCPE design method, the signal length required to satisfy $\text{rank}\ {H_{L}^{hyb}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}=mL$ is $T_{0}+\sum\limits_{i=\bar{p}}^{p}{{T_{i}}}\geq mL+(p-\bar{p}+1)(L-1)$.
Additionally, all three design methods ensure that each sub-signal is non-persistently exciting, i.e., $\text{rank}\ {H_{L}}({z_{i}^{[0,T_{i}-1]}})<mL$.
Proof. 1). The design choices ensure that the $p$ signals collectively construct the desired matrix while satisfying the length constraint $T_{i}<(m+1)L$ if $p>1$. Therefore,
$\text{rank}\ {H_{L}}({z_{i}^{[0,T_{i}-1]}})<mL.$
2). Since $\text{rank}\ \sum^{p}_{i=1}\alpha_{i}z_{i}\left(k\right)=1$ for $k\in\{L-1,2L-1,....,mL-1\}$, and $\sum^{p}_{i=1}\alpha_{i}z_{i}\left(k\right)=0$ otherwise, we have
$\text{rank}\ {H_{L}^{cum}(\{z^{[kL,(k+2)(L-1)]}_{i}\}_{i=1}^{p})}=L$
for all $k\text{}\in\left\{0,1,...,m-1\right\}$. At this point, $H_{L}^{cum}(\{z^{[kL,(k+2)(L-1)]}_{i}\}_{i=1}^{p})$} takes the form of a lower triangular matrix with nonzero diagonal elements, i.e.,
${H_{L}^{cum}(\{z^{[kL,(k+2)(L-1)]}_{i}\}_{i=1}^{p})}$
$=\begin{bmatrix}0&&&z^{cum}_{(k+1)(L-1)}\\ &&\iddots&z^{cum}_{k(L-1)+L}\\ &z^{cum}_{(k+1)(L-1)}&\iddots&\vdots\\ z^{cum}_{(k+1)(L-1)}&z^{cum}_{k(L-1)+L}&\cdots&z^{cum}_{(k+2)(L-1)}\end{bmatrix}$,
where $z^{cum}_{k}=\sum^{p}_{i=1}\alpha_{i}z_{i}\left(k\right)$.
Since $\sum^{p}_{i=1}\alpha_{i}\left[z_{i}\begin{matrix}\left(L-1\right)&z_{i}\left(2L-1\right)&\cdots&z_{i}\left(mL-1\right)\end{matrix}\right]$ has full rank, the choice of $z_{i}(k)$, $k\notin\{L-1,2L-1,....,mL-1\}$ at each time step increases the rank of the cumulative Hankel matrix $H_{L}^{cum}(\{z^{[0,T_{0}-1]}_{i}\}_{i=1}^{p})$} until it becomes full rank at $T_{0}=(m+1)L$.
Next, we demonstrate that designing $z_{i}(k)$ for k>mL1k>mL-1 according to the proposed approach ensures that each sub-Hankel matrix does not satisfy $mL$-order persistently exciting. Specifically, since $Z_{i}^{0}$ is full rank for $i=1,\dots,p$, there always exists a design law as described in (6). Consequently, we have
$z^{[k,k+L-1]}_{i}\in\text{im}\left[\begin{matrix}Z^{k-mL+1}_{i}\\ Z^{k-mL+2}_{i}\\ \vdots\\ Z^{k-(m-1)L}_{i}\end{matrix}\right],k\geqslant mL-1$
for all $i\in\{1,\dots,p\}$. This result implies that the rank of the Hankel matrix $H_{L}(z_{i}^{[0,T_{0}-1]})$ cannot increase beyond the $mL$-th column. Therefore,
$\text{rank}\ {H_{L}}({z_{i}^{[0,{T_{0}}-1]}})<mL.$
3). Building upon the proof procedures for signal design under the MCPE and CCPE conditions, the HCPE condition can be similarly achieved. Specifically, the matrix $H_{L}^{hyb}(\{z^{[0,T_{i}-1]}i\}{i=1}^{p})$ attains full row rank if $T_{0}+\sum_{j=\bar{p}+1}^{p}T_{j}\geq(m+1)L$, and each signal exhibits non-$mL$-order persistently exciting. $\blacksquare$
Remark 4.
Recent finite-time experiment design methods (e.g., Venkatasubramanian et al. (2025)) optimize a single exploration trajectory, typically harmonic or multisine inputs, to minimize parameter uncertainty and guarantee a priori performance bounds after identification. In contrast, the goal of our experiment design is not to minimize estimation error for a pre-specified model, but rather to ensure that the data collected across possibly multiple trials satisfy the proposed CPE conditions, which are sufficient for data-driven controller synthesis. This is crucial in settings where a single long PE trajectory cannot be executed due to safety or operational constraints. Moreover, our approach improves the numerical conditioning of the resulting Hankel matrix via weighting, thereby reducing the number of experimental repetitions needed to achieve data informativity. Hence, while targeted exploration focuses on minimizing model uncertainty for robust performance, our design explicitly guarantees the feasibility of direct data-driven control synthesis from multiple finite trajectories, even under limited excitation budgets.
Remark 5.
According to Theorem 2, all three conditions can be satisfied within the minimum number of time steps. Specifically, if we fix $H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})$}, $H_{L}^{cum}(\{z^{[0,T_{0}-1]}_{i}\}_{i=1}^{p})$}, and $H_{L}^{hyb}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})$} as standard square matrices, the experimental design method outlined in Section 4 can achieve the required conditions. In this scenario, the eigenvalues of the three composite matrices depend solely on the elements along the diagonal of each $L$ column. Furthermore, we note that the experimental design methods for all three conditions do not require knowledge of future information, allowing them to be implemented either offline or online. This enables real-time adjustment of control signals based on system state constraints while satisfying design requirements, thereby preventing extreme state divergence.
5 Application in Distributed Adaptive Identification
In this section, we explore the shift from single trajectories for identification to distributed identification using multiple trajectories that satisfy any CPE condition. Both single and distributed identifiers, our approach is direct, online and adaptive. Additionally, we highlight that the devised identification technique remains unaffected by the signal’s upper limit, making it highly robust as well.
Before proceeding, we establish exponential stability conditions for the autonomous system
$x(k+1)=(I_{n}-z(k)z^{T}(k))x(k),$ (7)
where $x(k)\in\mathbb{R}^{n}$, $z(k)\in\mathbb{R}^{n\times m}$, at the equilibrium point $x_{e}=0$. This system is a simplified error expression for the subsequent single-trajectory based identifier and multi-trajectory based distributed identifier. Moreover, it forms the theoretical foundation for analyzing the exponential convergence properties of the identifiers.
Lemma 3.
For the system (7), the equilibrium state $x_{e}=0$ is exponentially stable (e.s.) if the following conditions are satisfied for all $k\geq 0$:
  • i).
    The signal $z(k)$ is bounded, i.e. there exists a constant $z_{m}>0$ such that $z(k)||<{z_{m}}$;
  • ii).
    $2I-{z^{T}}(k)z(k)>0$;
  • iii).
    $z_{[k,k+l-1]}$ is PE of order $1$ for a fixed constant $l\geq 1$.
Proof: See the Appendix. $\blacksquare$
5.1 Adaptive Identifier for a Single Control Sequence Satisfying PE Condition
In this subsection, we devise an adaptive identifier for a single trajectory sequence within the framework established by Lemma 3, using the control sequence necessitating persistent excitation. In the subsequent subsections, we extend the results to distributed adaptive identifiers for multi-trajectory sequences, where each of the control sequence can come from either smooth or low-rank controller without having persistently exciting.
To facilitate system model identification, we transform the LTI system (1) into
$x(k+1)=r^{T}(k)\theta,$ (8)
where $r(k)=I_{n}\otimes\begin{bmatrix}{x(k)}\\ {u(k)}\end{bmatrix}\in{\mathbb{R}^{n(n+m)\times{n}}}$, $\theta=\begin{bmatrix}A&B\end{bmatrix}_{vec}\in{\mathbb{R}^{n(n+m)}}$.
For systems represented by (8), a discrete-time adaptive identifier is designed as
$\theta(k+1)=\theta(k)+\alpha r(k){\left({{r^{T}}(k)r(k)+\xi I}\right)^{-1}}\varepsilon(k),$ (9)
where $\theta(k)$ is the identification vector at moment $k$, $\alpha>0$ denotes the adaptive gain, $\xi>0$ is an arbitrary positive constant, and $\varepsilon(k)={x(k+1)-{r^{T}}(k)\theta(k)}$.
Then, the dynamics of the identification error $\tilde{\theta}(k)=\theta(k)-\theta$ of (9) can be expressed as
$\tilde{\theta}(k+1)=(I-\alpha r(k){\left({{r^{T}}(k)r(k)+\xi I}\right)^{-1}}{r^{T}}(k))\tilde{\theta}(k).$ (10)
Since ${r^{T}}(k)r(k)+\xi I$} is an invertible symmetric matrix, (10) possesses the same form as (7). We can use Lemma 3 to establish the exponential convergence condition of the identifier (9), as shown in the following theorem.
Theorem 3.
For the identifier (9), $\theta(k)\rightarrow\theta$ exponentially fast if the following conditions hold for all $k\geq 0$:
  • i).
    $r(k)$ is bounded;
  • ii).
    $0<\alpha\leq 2$;
  • iii).
    $r_{[k,k+l-1]}$ satisfies PE condition of order $1$ with a fixed constant $l\geq 1$.
Proof: Utilizing the Woodbury Matrix Identity, we obtain $r{({r^{T}}r+\xi I)^{-1}}{r^{T}}=I-{(I+\frac{1}{\zeta}r{r^{T}})^{-1}}<I$. Consequently,
$2I-\alpha r(k){\left({{r^{T}}(k)r(k)+\xi I}\right)^{-1}}{r^{T}}(k)>0$ (11)
when $0<\alpha\leq 2$, thereby fulfilling condition ii) of Lemma 3. The remaining assumptions in Theorem 3 correspond to conditions i) and iii) in Lemma 3, respectively. Hence, $\tilde{\theta}=0$ is exponentially stable, implying that $\theta(k)$ converges exponentially fast to $\theta$. $\blacksquare$
Note that for the systems described by (1) or (8), when a set of input-state data can uniquely determine $A$ and $B$, it is typically necessary for the input sequence to constitute a persistently exciting sequence of order $n+1$. If the excitation condition for $u_{[k,k+l-1]}$ holds for all $k\geq 0$ with a fixed constant $l\geq 1$, then $\begin{bmatrix}{{H_{1}}({x_{[k,k+l-1]}})}\\ {{H_{1}}({u_{[k,k+l-1]}})}\end{bmatrix}$ achieves full row rank. Consequently, ${H_{1}}({r_{[k,k+l-1]}})$ also attains full row rank, meeting the excitation requirement outlined in Theorem 3. Therefore, the following corollary ensues.
Corollary 1.
For the identifier (9), $\theta(k)\rightarrow\theta$ exponentially fast if the following conditions hold for all $k\geq 0$:
  • i).
    $x(k)$ and $u(k)$ are bounded;
  • ii).
    $0<\alpha\leq 2$;
  • iii).
    $u_{[k,k+l-1]}$ satisfies PE condition of order $n+1$ with a fixed constant $l\geq 1$.
5.2 Distributed Adaptive Identifiers for Multiple Control Sequences Satisfying CPE Conditions
In this subsection, we will leverage the CPE condition to formulate distributed adaptive identifiers with smooth input signals within the state-space setting, elucidating their conditions for exponential stability. Each sequence of trajectories used need not satisfy the identifiability condition, and can also enables the synthesis of control inputs for the desired system through the stabilizing controllers of multiple subsystems without necessitating perturbation experiments.
Consider a homogeneous multi-agent system (MAS) with $N$ agents, the dynamics of each agent are given by
$x_{i}(k+1)=r_{i}^{T}(k)\theta,\ \ i=1,2,...,N$ (12)
where $r_{i}(k)=I\otimes\begin{bmatrix}{x_{i}(k)}\\ {u_{i}(k)}\end{bmatrix}\in{\mathbb{R}^{n(n+m)\times{n}}}$, $\theta=\begin{bmatrix}A&B\end{bmatrix}_{vec}\in{\mathbb{R}^{n(n+m)}}$.
Using the rank condition under CPE provided in Lemma 1, there is a centralized non-adaptive identification scheme similar to (4). Taking the CCPE condition as an example, one obtains
${H_{1}^{c}({x_{1[1,{T_{0}}]}^{p}})}=\begin{bmatrix}A&B\end{bmatrix}\begin{bmatrix}{H_{1}^{c}({x_{1[0,{T_{0}-1}]}^{p}})}\\ {H_{1}^{c}({u_{1[0,{T_{0}-1}]}^{p}})}\end{bmatrix}$ (13)
If $T_{0}=T_{1}=T_{2}=...=T_{p}$ and $\{u_{i[0,{T_{i}-1}]}\}_{i=1}^{p}$ are CCPE of order $n+1$, then $\begin{bmatrix}{H_{1}^{c}({x_{1[0,{T_{0}-1}]}^{p}})}\\ {H_{1}^{c}({u_{1[0,{T_{0}-1}]}^{p}})}\end{bmatrix}$ has full row rank. Hence, we have
$\begin{bmatrix}A&B\end{bmatrix}={H_{1}^{c}({x_{1[1,{T_{0}}]}^{p}})}{\begin{bmatrix}{H_{1}^{c}({x_{1[0,{T_{0}-1}]}^{p}})}\\ {H_{1}^{c}({u_{1[0,{T_{0}-1}]}^{p}})}\end{bmatrix}^{\dagger}}$, (14)
where $\dagger$} denotes the right inverse.
However, this centralized non-adaptive identification method is complex and poses a higher risk of being vulnerable to covert and stealthy attacks. Additionally, it exhibits poor identification capability in the presence of changing parameters. Therefore, in the following, we delve into the design of distributed adaptive identifiers to tackle these challenges.
Based on the previous identifier (9) for a single trajectory, the distributed identifier for the system (12) is designed as
${\theta_{i}}(k+1)={\theta_{i}}(k)+$ $\alpha{r_{i}}(k){\left({{r_{i}}^{T}(k){r_{i}}(k)+\xi I}\right)^{-1}}{\varepsilon_{i}}(k)$
$~~~~~~-\gamma\sum\limits_{j\in{\mathcal{N}_{i}}}{\left({{\theta_{i}}(k)-{\theta_{j}}(k)}\right)},$ (15)
where $\theta_{i}(k)$ is the identification vector of $i$-th agent at moment $k$, $\alpha>0$ and $\gamma>0$ are constants to be designed, $\xi>0$ is an arbitrary positive constant, and $\varepsilon_{i}(k)={x_{i}(k+1)-{r_{i}^{T}}(k)\theta_{i}(k)}$.
Let ${{\tilde{\theta}}_{i}}(k)={\theta_{i}}(k)-\theta$ and ${\tilde{\theta}_{c}}(k)=\begin{bmatrix}{\tilde{\theta}_{1}^{T}(k)},{\tilde{\theta}_{2}^{T}(k)},...,{\tilde{\theta}_{N}^{T}(k)}\end{bmatrix}^{T}$ denote the identification error of the $i$-th identifier and the compact form of the errors, respectively. Then, the error dynamics can be expressed as
$\tilde{\theta}_{c}(k+1)=\left({I-\alpha R(k)-\gamma{\mathcal{L}}\otimes I}\right)\tilde{\theta}_{c}(k),$ (16)
where $R(k)=\text{diag}(\bar{r}_{1}(k),...,\bar{r}_{N}(k))$ and $\bar{r}_{i}(k)=r_{i}(k){\left({{r_{i}^{T}}(k)r_{i}(k)+\xi I}\right)^{-1}}{r_{i}^{T}}(k)$.
Theorem 4.
Suppose that the communication network $\mathcal{G}$ is undirected. Then, $\theta_{i}(k)\rightarrow\theta$, i=1,2,,Ni=1,2,...,N exponentially fast if the following conditions hold for all $k\geq 0$:
  • i).
    The input-state data of the MAS is bounded;
  • ii).
    $0<\alpha\leq 2$ and $0<\gamma<\frac{1}{{{\lambda_{\max}}(\mathcal{L})}}$;
  • iii).
    The inputs $\{u_{i[k,{k+l-1}]}\}_{i=1}^{N}$ are any type of CPE of order $n+1$ for a fixed constant $l\geq 1$.
Proof: By definition, we know that $\alpha R(k)+\gamma{\mathcal{L}}\otimes I$ is a semi-positive definite matrix, so there exists a matrix $\Phi(k)$ such that $\Phi(k)\Phi^{T}(k)=\alpha R(k)+\gamma{\mathcal{L}}\otimes I$. Then, (16) can be written as
$\tilde{\theta}_{c}(k+1)=\left({I-\Phi(k)\Phi^{T}(k)}\right)\tilde{\theta}_{c}(k).$ (17)
We first prove that $\Phi_{[k,k+l-1]}$ is PE of order $1$ for a fixed constant $l\geq 1$ when $\{u_{i[k,{k+l-1}]}\}_{i=1}^{N}$ are MCPE of order $n+1$. The problem can be translated into proving that
$\sum\limits_{t=0}^{l-1}{\Phi(k+t){\Phi^{T}}(k+t)}=l\gamma\mathcal{L}\otimes I+\alpha\sum\limits_{t=0}^{l-1}{R(k+t)}>0.$
The preceding inequality can be equivalently expressed as $\ker\mathcal{L}\otimes I\cap\ker\sum\limits_{t=0}^{l-1}{R(k+t)}=\{0\}$. In graph theory, for an undirected and connected communication network $\mathcal{G}$, $\mathcal{L}\otimes I$ possesses $(n+m)n$ zero eigenvalues. The elements in $\ker\mathcal{L}\otimes I$ can be represented as $(\textbf{1}_{(n+m)n}\otimes I_{(n+m)n})\eta$, where $\eta\in\mathbb{R}^{(n+m)n}$. Consequently, we obtain
${\eta^{T}}{({{\bf{1}}_{(n+m)n}}\otimes{I_{(n+m)n}})^{T}}\sum\limits_{t=0}^{l-1}{R(k+t)}({{\bf{1}}_{(n+m)n}}\otimes{I_{(n+m)n}})\eta$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~={\eta^{T}}\sum\limits_{t=0}^{l-1}{\sum\limits_{i=1}^{N}{{{\bar{r}}_{i}}(k+t)}}\eta.$ (18)
If $\{u_{i[k,{k+l-1}]}\}_{i=1}^{N}$ satisfy the MCPE condition, it can be readily inferred that $\sum\limits_{t=0}^{l-1}{\sum\limits_{i=1}^{N}{{{\bar{r}}_{i}}(k+t)}}>0$. Thus, (5.2) equals 0 if and only if $\eta$ is the zero vector, implying that $\ker\mathcal{L}\otimes I\cap\ker\sum\limits_{t=0}^{l-1}{R(k+t)}=\{0\}$. Moreover, according to Lemma 1, if $\{u_{i[k,{k+l-1}]}\}_{i=1}^{N}$ satisfy the CCPE or HCPE condition, they must also satisfy the MCPE condition. Therefore, $\Phi_{[k,k+l-1]}$ are PE of order $1$ for all types of CPE of order $n+1$.
Next we prove that $2I-\Phi^{T}(k)\Phi(k)>0$. Let $\alpha R(k)=\phi_{1}(k)\phi_{1}^{T}(k)$ and $\gamma{\mathcal{L}}\otimes I=\phi_{2}(k)\phi_{2}^{T}(k)$, thus we have $\Phi(k)=\begin{bmatrix}{\phi_{1}(k)}&{\phi_{2}(k)}\end{bmatrix}$. Then, the inequality $2I-\Phi^{T}(k)\Phi(k)>0$ is equivalent to
$\begin{bmatrix}{2I-\phi_{1}^{T}(k){\phi_{1}}(k)}&-{\phi_{1}^{T}(k){\phi_{2}}(k)}\\ -{\phi_{2}^{T}(k){\phi_{1}}(k)}&{2I-\phi_{2}^{T}(k){\phi_{2}}(k)}\end{bmatrix}>0.$ (19)
By (11), if $0<\alpha\leq 2$, then $2I-\phi_{1}^{T}(k){\phi_{1}}(k)>0$. Thus, the Schur complement of (19) is of the form
$2I-\phi_{2}^{T}(k){\phi_{2}}(k)-\phi_{2}^{T}(k)M_{2}(k){\phi_{2}}(k)>0,$
where $M_{2}(k)={\phi_{1}}(k){\left({2I-\phi_{1}^{T}(k){\phi_{1}}(k)}\right)^{-1}}\phi_{1}^{T}(k)$. According to the Woodbury Matrix Identity, ${M_{2}}(k)=I-{\left({I+\frac{1}{2}{\phi_{1}}(k)\phi_{1}^{T}(k)}\right)^{-1}}<I$, which implies that
$2I-\phi_{2}^{T}(k){\phi_{2}}(k)-\phi_{2}^{T}(k){M_{2}}(k){\phi_{2}}(k)>2I-2\phi_{2}^{T}(k){\phi_{2}}(k).$
Hence, (19) holds if $0<\gamma<\frac{1}{{{\lambda_{\max}}(\mathcal{L})}}$. In conclusion, the conditions given in Theorem 4 ensure that the conditions in Lemma 3 hold, so we have $\theta_{i}(k)\rightarrow\theta$, i=1,2,,Ni=1,2,...,N exponentially fast. $\blacksquare$
Remark 6.
Both the single identifier (9) and the distributed identifiers (5.2) rely solely on specific constants for selecting parameters within their exponentially stabilized convergence conditions. They maintain robustness to trajectory data, imposing no stringent requirements on the exact values of upper bounds, as long as signal trajectories remain bounded.
6 Illustrative Examples of Applications
6.1 Distributed Identification from Multiple Control Inputs
In this subsection, we apply the distributed adaptive identification scheme of Section 5 to the voltage converter systems. Consider a homogeneous multi-agent systems comprising five agents, and the communication network $\mathcal{G}$ is shown in Fig. 2. By discretizing the voltage converter system described in Tan and Hoo (2015) with a sampling time of 0.1 seconds, we derive the state-space representation for each agent as
$x_{i}(k+1)=\begin{bmatrix}{1.0000}&{-0.0500}\\ {0.0004}&{\ \ 0.9998}\end{bmatrix}x_{i}(k)+\begin{bmatrix}{0.0125}\\ {0.0000}\end{bmatrix}u_{i}(k),$
where $i\in\{1,2,3,4,5\}$. Transforming the system into the form of (12) results in $\theta=\begin{bmatrix}1.0000&{-0.0500}&{0.0125}&{0.0004}&{0.9998}&0.0000\end{bmatrix}^{T}$, where $\theta$ represents the parameters subject to identification.
Figure 2: Network topology composed of all agents.
We choose $\xi=2$, $\alpha=1$ and $\gamma=0.25$, satisfying condition ii) in Theorem 4. A comparison is conducted between the distributed identification scheme delineated in Section 5 and the single-system identification scheme. In the single-system identification scheme, we consider cases with and without random noise. The addition of random noise ensures that each subsystem satisfies the persistently exciting condition in Theorem 3 and thus meets the identifiable condition. Conversely, the case without noise is employed to demonstrate the distributed identification scheme’s resilience against covert and stealthy attacks.
Figure 3: Distributed adaptive identification from multiple sequences of feedback control inputs.
Figure 4: Single-system identification from a single input sequence with random noise.
Figure 5: Single-system identification from a single input sequence without random noise.
Fig. 3 depicts the distributed adaptive identification results with the collaborative efforts of five agents. It is evident that each agent is able to approximate the desired identification result with a relatively smooth curve when the CPE condition is satisfied. This smoothness arises from the absence of additional noise introduced by the agents’ controllers, allowing the multi-agent system to perform the identification through interaction using the smooth system trajectories.
In contrast, the identification result of a single system, as depicted in Fig. 4, confirm the identification ability of the identifier (9) under persistently exciting condition (see Corollary 1). However, due to the requirement for sufficiently stimulating input data, the random noise is injected to the controller, resulting in significant fluctuations during the single-system identification process. The successful identification also indicates a lack of capability to withstand covert, stealth attacks, which is one of the motivations behind our research.
Fig. 5 displays the identification results of a single system without the addition of random noise. It is evident that due to the absence of identifiable conditions, the identifier converges to undesirable results in a short period. This demonstrates the challenge of achieving convergence to the target model using single-input trajectory single-system identification under the influence of bounded feedback controllers without satisfying sustained excitation conditions, further affirming the superiority of the distributed adaptive identification proposed in this paper under CPE conditions. Additionally, it illustrates the effectiveness of distributed identification methods in reducing the risk posed by attacks that require precise modeling.
6.2 Least-squares Identification from Multiple Control Inputs
In this subsection, we apply the LS identification method to a batch reactor system, previously considered in De Persis and Tesi (2020). The system dynamics are described by
$x_{i}(k+1)$ $=\begin{bmatrix}{1.178}&{0.001}&{0.511}&{-0.403}\\ {-0.051}&{0.661}&{-0.011}&{0.061}\\ {0.076}&{0.335}&{0.560}&{0.382}\\ 0&{0.335}&{0.089}&{0.849}\end{bmatrix}x_{i}(k)$
$~~~~~~~~~~~~~~~~~~~~~+\begin{bmatrix}{0.004}&{-0.087}\\ {0.467}&{0.001}\\ {0.213}&{-0.235}\\ {0.213}&{-0.016}\end{bmatrix}u_{i}(k)+w_{i}(k),$ (20)
where $w_{i}(k)\in{\mathbb{R}^{n}}$ represents the process noise. It is assumed that the process noises is i.i.d. Gaussian, i.e., $w_{i}(k)\sim\mathcal{N}(0,\sigma_{w}^{2}I)$.
This system has a state order of $n=4$, input order of $m=2$, and is open-loop unstable.
For each agent, the input/state data is organized as
$U_{i}=\left[\begin{matrix}u_{i}\left(0\right)&u_{i}\left(1\right)&\cdots&u_{i}\left(T_{i}-1\right)\end{matrix}\right],$
$X_{i}^{-}=\left[\begin{matrix}x_{i}\left(0\right)&x_{i}\left(1\right)&\cdots&x_{i}\left(T_{i}-1\right)\end{matrix}\right],$
$X_{i}^{+}=\left[\begin{matrix}x_{i}\left(1\right)&x_{i}\left(2\right)&\cdots&x_{i}\left(T_{i}\right)\end{matrix}\right],$
and the process noise data as
$W_{i}=\left[\begin{matrix}w_{i}\left(0\right)&w_{i}\left(1\right)&\cdots&w_{i}\left(T_{i}-1\right)\end{matrix}\right],$
$W_{i}^{n}=\left[\begin{matrix}w_{i}\left(0\right)&w_{i}\left(1\right)&\cdots&w_{i}\left(T_{i}-1-n\right)\end{matrix}\right].$
The data combinations in Definitions 2-4 can be freely selected according to the length of the input data of the $p$ agents. For example, consider the CCPE condition. If $T_{1}=T_{2}=...=T_{p}=T_{0}$, the input data can be combined as
$H^{cum}_{L}\left(\{U_{i}\}^{p}_{i=1}\right)=\sum\limits^{p}_{i=1}{{\alpha_{i}}{H_{L}}(U_{i})}.$
From the system dynamics (6.2), we obtain
$H^{cum}_{1}\left(\{X^{+}_{i}\}^{p}_{i=1}\right)$
$=\begin{bmatrix}A&B\end{bmatrix}\left[\begin{matrix}H^{cum}_{1}\left(\{X^{-}_{i}\}^{p}_{i=1}\right)\\ H^{cum}_{1}\left(\{U_{i}\}^{p}_{i=1}\right)\end{matrix}\right]+H^{cum}_{1}\left(\{W_{i}\}^{p}_{i=1}\right).$
Define $G=\begin{bmatrix}A&B\end{bmatrix}$. The least squares problem is formulated as
$\text{}\min_{G}\left\|H^{cum}_{1}\left(\{X^{+}_{i}\}^{p}_{i=1}\right)-G\left[\begin{matrix}H^{cum}_{1}\left(\{X^{-}_{i}\}^{p}_{i=1}\right)\\ H^{cum}_{1}\left(\{U_{i}\}^{p}_{i=1}\right)\end{matrix}\right]\right\|$
The solution, unique due to Lemma 1, is given by
$\hat{G}={H^{cum}_{1}(\{X_{i}^{+}\}^{p}_{i=1})}{\begin{bmatrix}{H^{cum}_{1}(\{X_{i}^{-}\}^{p}_{i=1})}\\ {H^{cum}_{1}(\{U_{i}\}^{p}_{i=1})}\end{bmatrix}^{{\dagger}}}$. (21)
In this example, we employ the CCPE, MCPE, and HCPE conditions for LS identification under three different data settings: same-dimensional, different-dimensional, and partially same-dimensional data, respectively. According to (21), $\hat{G}$ is identifiable if the matrix $\begin{bmatrix}{H^{cum}_{1}(\{X_{i}^{-}\}^{p}_{i=1})}\\ {H^{cum}_{1}(\{U_{i}\}^{p}_{i=1})}\end{bmatrix}$ has full row rank. Lemma 1 indicates that this requirement holds as long as the order of the CPE satisfies $L\geq n+1$. For systems with order $n=4$, this implies that the minimum required order of the CPE is $5$. The choice of $L$ involves a trade-off between data length, numerical conditioning, and computational cost. A larger $L$ can improve subspace separation but requires longer experiments and heavier computation, and may worsen conditioning under noise. In practice, it is recommended to select the smallest $L$ that meets the theoretical rank condition while ensuring acceptable numerical properties and manageable experiment duration. To synthesize the necessary CPE conditions of order $5$, we utilize 10 independently input trajectories ($p=10$), each of length $T_{i}$ samples.
Figure 6: Identification errors under different settings. (a)-(c) Identification results using the MCPE, CCPE, and HCPE conditions, respectively, across different noise levels and synthetic matrix singular value settings. (d) Identification results for the MCPE condition under various weighting factor configurations.
Figure 7: Control input sequences generated under the MCPE, CCPE, and HCPE conditions, respectively, along with the total input lengths and the corresponding Hankel matrix column sizes.
For the MCPE condition, the data sets can have varying dimensions but must satisfy $\sum_{i=1}^{p}T_{i}\geq 50$ (see Theorem 2). Control sequence lengths are randomly generated within the range [5, 14). For the CCPE condition, the ten data sets must have identical dimensions and satisfy $T_{0}\geq 14$ (see Theorem 2), we set $T_{0}=25$. This larger choice is deliberate and provides a practical safety margin against measurement noise, finite-precision arithmetic, and unmodeled dynamics. It also increases the number of columns $T_{0}-L+1$, which makes the ensuing least-squares and pseudoinverse computations more overdetermined and numerically stable. For the HCPE condition, partial dimensional consistency is allowed $\bar{p}=3$, with the condition $T_{0}+\sum_{i=\bar{p}}^{p}T_{i}\geq 42$ (see Theorem 2). Let $T_{0}=10$, the remaining controller lengths are randomly generated within [5, 14).
The input sequences for all three conditions are designed using the method in Section 4. The initial state values are randomly generated within the range [-1, 1]. Additionally, the standard deviation of the process noise $\sigma_{w}$ is varied from 0 to $0.1$. The weighting factors for each CPE condition are fixed as $\alpha_{i}=1$ for $i=1,2,\dots,10$. we solve for $\hat{G}$ in (21) using the pinv function from the standard MATLAB Linear Algebra Toolbox. The identification errors under various conditions are shown in Fig. 6(a)-(c). The results validate the effectiveness of the three synthesized conditions for least-squares identification.
To highlight the significance of weighting factors, we examined scenarios where $\alpha_{1}$ and $\alpha_{2}-\alpha_{10}$ varied within the range (0, 30] under the MCPE condition. The results, corresponding to a noise level of $\sigma_{w}=0.05$, are presented in Fig. 6(d). The red line represents the case where $\alpha_{1}$ varies while $\alpha_{2}-\alpha_{10}$ remain fixed at 1, whereas the blue line represents the case where $\alpha_{1}$ is fixed at 1 and $\alpha_{2}-\alpha_{10}$ vary synchronously. The findings indicate that significant disparities in signal magnitudes under the MCPE condition can degrade discrimination performance. Conversely, an appropriate selection of weighting factors effectively mitigates identification errors.
6.3 State Feedback Controller Design Using Parametric Modeling Approach
In this subsection, we apply the data-driven state feedback design method described in De Persis and Tesi (2020) to the system defined by (6.2) using the three CPE conditions. Specifically, taking the CCPE condition as an example. Utilizing the approach in (De Persis and Tesi, 2020, Th. 3), the design challenge for stabilizing controllers is transformed into an LMI problem for any matirx $Q$:
$\begin{bmatrix}{{H_{1}^{cum}(\{{x_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})}Q}&{{H_{1}^{cum}(\{{x_{i}^{[1,{T_{0}}]}\}_{i=1}^{p}})}Q}\\ {{{\left({{H_{1}^{cum}(\{{x_{i}^{[1,{T_{0}}]}\}_{i=1}^{p}})}Q}\right)}^{T}}}&{{H_{1}^{cum}(\{{x_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})}Q}\end{bmatrix}>0,$ (22)
and the feedback control gain can be obtained by
$K_{cum}=$ ${H_{1}^{cum}(\{{u_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})}Q$
$~~~~~~~~~~~\times{\left({{H_{1}^{cum}(\{{x_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})}Q}\right)^{-1}}.$ (23)
Similar to LS identification, (22) and (6.3) are feasible if and only if the corresponding CPE conditions satisfy at least the $n+1$ order. For this example, we use the control signals generated in the previous subsection as a priori control inputs to generate corresponding input-state trajectories for the three CPE conditions of order $5$. Fig. 7 presents the control input sequences generated under the MCPE, CCPE, and HCPE conditions, respectively. Each subplot corresponds to one excitation strategy and depicts the concatenated input signals of the first control component across ten trajectory segments.
In the case of the CCPE condition, the collected input-state trajectories are synthesized into $H_{1}^{cum}(\{{u_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})$}, $H_{1}^{cum}(\{{x_{i}^{[0,{T_{0}-1}]}\}_{i=1}^{p}})$}, and $H_{1}^{cum}(\{{x_{i}^{[1,{T_{0}}]}\}_{i=1}^{p}})$}. These matrices are then processed using MATLAB’s LMI toolbox Duan and Yu (2013) to compute the control gain $K_{cum}$ using (22) and (6.3):
${K_{cum}}=\begin{bmatrix}{0.8480}&{-1.4693}&{0.1463}&{-1.5678}\\ {4.0457}&{0.0394}&{3.2812}&{-1.4414}\end{bmatrix}$.
It can be verified that the closed-loop matrix $A+BK_{cum}$ is stable with a spectral radius of $0.5922$.
For the MCPE condition, the collected input-state trajectories are synthesized into $H_{1}^{mos}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})$}, $H_{1}^{mos}(\{{x_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})$}, and $H_{1}^{mos}(\{{x_{i}^{[1,{T_{i}}]}\}_{i=1}^{p}})$}, and these matrices replace the corresponding matrices in (22) and (6.3). Solving the resulting LMI yields the control gain $K_{mos}$:
${K_{mos}}=\begin{bmatrix}{0.8218}&{-1.4683}&{0.1403}&{-1.5406}\\ {3.9691}&{0.0432}&{3.2652}&{-1.3369}\end{bmatrix}$.
It can be verified that the closed-loop matrix $A+BK_{mos}$ is stable with a spectral radius of $0.5921$.
Finally for the HCPE condition, the resulting input-state trajectories are organized into $H_{1}^{hyb}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})$}, $H_{1}^{hyb}(\{{x_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})$}, and $H_{1}^{hyb}(\{{x_{i}^{[1,{T_{i}}]}\}_{i=1}^{p}})$}, replacing the corresponding matrices in (22) and (6.3). Solving yields a control gain $K_{hyb}$:
${K_{hyb}}=\begin{bmatrix}{0.8416}&{-1.4690}&{0.1450}&{-1.5594}\\ {4.0254}&{0.0399}&{3.2761}&{-1.4282}\end{bmatrix}$.
It can be verified that the closed-loop matrix $A+BK_{hyb}$ is stable with a spectral radius of $0.5942$.
6.4 Data-Driven MPC Under Three CPE Conditions
In this subsection, we apply the data-driven MPC method to a batch reactor system considered in the previous subsection.
It is feasible to directly employ Lemma 2 for the MPC problem:
[…]
where $N$ is the predicted step size, $u^{[k-(n-1),k]}$ and $x^{[k-(n-1),k]}$ are the k(n1)k-(n-1) moment to $k$ moment individual input-state trajectories, respectively, which are taken as the initial values at moment $k$. ${\bar{u}}^{[1,N]}$} and ${\bar{x}}^{[1,N]}$}} are the trajectories to be optimized. The sets $\mathbb{U}$ and $\mathbb{X}$ describe feasible inputs and states, respectively. When the cost function of MPC is a quadratic stage cost, i.e.,
${J\left({{{\bar{x}}_{m}},{{\bar{u}}_{m}}}\right)}=({{{\bar{x}}_{m}(k)}}-x^{*})^{T}Q({{{\bar{x}}_{m}(k)}}-x^{*})+{{{\bar{u}}_{m}^{T}(k)}}R{{{\bar{u}}_{m}(k)}},$
where $Q$, $R>0$, the problem can be transformed into a standard quadratic programming (QP) problem.
The primary goal is to track the set point of the system $x^{*}=[0\ 0\ 0\ 0]^{T}$. We select a prediction horizon of $N=5$ and set the weight matrices in the cost function as $Q=3\cdot I_{n}$ and $R=10^{-2}\cdot I_{m}$. There are no constraints imposed on the upper and lower bounds of the control values.
We employed the proposed experimental design approach to collect the desired priori data satisfying three CPE conditions. For each condition, 10 sets of control sequences were collected, each with 30 steps. In Fig. 8, we present the closed-loop states obtained by implementing the data-driven MPC scheme under CCPE, MCPE, and HCPE conditions, respectively. Here, $x^{i}$, i=1,2,,ni=1,2,...,n represents the $i$th state component of the closed-loop system. The outcomes demonstrate that the three methods exhibited nearly consistent (i.e., optimal) performance.
Furthermore, we compared the optimization times of the three methods, as shown in Table 1. The MCPE method, which has the largest number of columns in the synthesis matrix and the highest computational complexity, required the longest computation time. Conversely, the CCPE method had the shortest computation time, while the HCPE method was intermediate. These results are consistent with our anticipated outcomes.
Figure 8: Closed-loop states for data-driven MPC. (a). First state component. (b). Second state component. (c). Third state component. (d). Fourth state component.
Table 1: Running Times of Different Conditions
Condition Running Time (s)
CCPE 0.1919
MCPE 3.1657
HCPE 1.8771
6.5 Complexity Discussion Between MCPE and PE Conditions
Concrete numeric example (heterogeneous short trajectories): Let $m=2$ and $L=5$. For a trajectory $u^{[0,T_{0}-1]}$, the Hankel matrix $H_{L}(u^{[0,T_{0}-1]})\in\mathbb{R}^{mL\times C_{0}}$ with column counts $C_{0}=T_{0}{-}L{+}1$. Recall that a single-trajectory PE condition requires
$T_{0}\ \geq\ (m+1)L-1\;=\;3\cdot 5-1\;=\;14.$
Suppose long experiments are infeasible, but we can repeatedly collect short trials. Take five short trajectories with lengths
$T_{1}=7,\;T_{2}=7,\;T_{3}=6,\;T_{4}=6,\;T_{5}=5.$
Individually none satisfies $T_{i}\geq 14$. The mosaic Hankel matrix ${H_{L}^{mos}(\{u^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}\in\mathbb{R}^{mL\times C}$ with column counts $C=\sum_{i=1}^{N}\big(T_{i}-L+1\big)$. Their available column counts are
$T_{1}{-}L{+}1$ $=3,\;T_{2}{-}L{+}1=3,\;T_{3}{-}L{+}1=2,$
$T_{4}{-}L{+}1=2,\;T_{5}{-}L{+}1=1,$
so the mosaic Hankel matrix has
$C=3+3+2+2+1=11\;\;\geq\;mL=10.$
Under mild genericity (inputs not persistently collinear and weights $\alpha_{i}$ chosen to avoid near-cancellations), this suffices to achieve $\mbox{rank}(H^{mos}_{L})=mL$. Hence the MCPE test can be satisfied by one unified rank check on a $(mL)\times C=10\times 11$ matrix, even though no single trajectory is long enough to satisfy PE.
Complexity accounting: Let $\text{flops}_{(\cdot)}$ denote the leading-order cost of a rank test via singular value decomposition (SVD).
For the PE condition, one rank test on $\mathbb{R}^{mL\times C_{0}}$ with $C_{0}=T_{0}{-}L{+}1\geq mL$. A typical cost model gives
$\text{flops}_{\text{PE}}\;\sim\;\mathcal{O}\big((mL)^{2}C_{0}\big).$
However, when a long trajectory is not available upfront, one often repeats experiments and re-checks rank $K$ times until Definition 1 is met, giving a cumulative
$\text{flops}_{\text{PE, repeated}}\;\sim\;\sum_{k=1}^{K}\mathcal{O}\big((mL)^{2}C_{k}\big),$ (24)
with added experimental overhead for each trial.
For the MCPE condition, a single rank test on $\mathbb{R}^{mL\times C}$ with $C=\sum_{i}(T_{i}{-}L{+}1)_{+}$:
$\text{flops}_{\text{MCPE}}\;\sim\;\mathcal{O}\big((mL)^{2}C\big).$ (25)
Comparing (24) and (25), repeated PE testing becomes more expensive once
$\sum_{k=1}^{K}C_{k}>C,$ (26)
that is, when the cumulative number of columns generated by repeated trials exceeds that of the single mosaic matrix. If each trial produces of PE condition on average $\bar{c}$ columns, the threshold occurs at
$K_{\mathrm{th}}=\Big\lceil\tfrac{C}{\bar{c}}\Big\rceil.$ (27)
In the previous example, if each failed PE attempt produces is $\bar{c}=C_{0}=10$ columns, then by (27) the threshold is $K_{\mathrm{th}}=\lceil 11/10\rceil=2$. Hence after two repeated PE trials the cumulative cost exceeds that of a single MCPE check.
This illustrates that, although MCPE involves a wider matrix than a one-shot PE check, it avoids repeated failed trials and preprocessing, thereby reducing the overall computational and experimental burden.
7 Conclution
In this paper, we have developed and analyzed three collectively persistently exciting (CPE) conditions, each incorporating weight factors, from the perspectives of same-dimensional, different-dimensional, and partially same-dimensional data. We argue that these three CPE conditions offer valuable insights for designing data-driven methodologies tailored to diverse multi-signal control scenarios, including those involving signals with insufficient informativeness. We have explored the interrelations between these conditions and examined their rank properties within the framework of linear time-invariant systems. All three conditions successfully extend Willems’ fundamental lemma. To address the challenges of insufficient information in signal segments, we have proposed open-loop experimental design methods tailored to each CPE condition, enabling the synthesis of the required excitation conditions both offline or online. Illustrative examples show that these conditions lead to satisfactory results in both model identification and the development of data-driven controllers.
Several topics for future research remain open. An interesting but challenging extension lies in devising data-driven control methods in a distributed framework, as opposed to a centralized one, while ensuring adherence to the three CPE conditions with multiple control signals. Furthermore, we note that weighting factors have an impact on the performance and effectiveness of data-driven methods. Therefore, how to systematically design optimal weighting factors for data-driven problems is another direction worth exploring. In addition, it is of interest to investigate how different experiment design methods influence parameter uncertainties and control performance, thereby integrating uncertainty-aware objectives into the proposed framework.
Appendix
A. Proof of Theorem 1
1). If $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are CCPE, then $T_{0}=T_{1}=T_{2}=...=T_{p}$ and ${H_{L}^{cum}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}{{H_{L}^{cum}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}}^{T}>0$, i.e.,
$\sum\limits_{i=1}^{p}{\left({{\alpha_{i}}{H_{L}}({z_{i}^{[0,{T_{i}}-1]}})\sum\limits_{i\neq j=1}^{p}{{\alpha_{j}}H_{L}^{T}({z_{j}^{[0,{T_{j}}-1]}})}}\right)}$
$~~~~~~~~~~+\underbrace{\alpha_{i}^{2}\sum\limits_{i=1}^{p}{{H_{L}}({z_{i}^{[0,{T_{i}}-1]}})H_{L}^{T}({z_{i}^{[0,{T_{i}}-1]}})}}_{M_{1}}>0.$ (28)
In order to establish the full row rank of $H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})$}, it is sufficient to show the positivity of $M_{1}$. Given $\sum\limits_{i=1}^{p}{{H_{L}}({z_{i}^{[0,{T_{i}}-1]}})H_{L}^{T}({z_{i}^{[0,{T_{i}}-1]}})}\geq 0$ and $\alpha_{i}\neq 0$, if there exists a non-zero vector $x$ such that $x\in\ker M_{1}$, then we must have $H_{L}^{T}({z_{1}^{[0,{T_{i}}-1]}})x=\cdots H_{L}^{T}({z_{p}^{[0,{T_{i}}-1]}})x=0$, implying
$x\in\ker\sum\limits_{i=1}^{p}{\left({{\alpha_{i}}{H_{L}}({z_{i}^{[0,{T_{i}}-1]}})\sum\limits_{i\neq j=1}^{p}{{\alpha_{j}}H_{L}^{T}({z_{j}^{[0,{T_{j}}-1]}})}}\right)}.$
This conclusion contradicts inequality (A. Proof of Theorem 1). Therefore $M_{1}>0$. Consequently, $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also MCPE.
Since ${H_{L}^{cum}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}={H_{L}^{cum}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{\bar{p}}})}+{H_{L}^{cum}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=\bar{p}+1}^{p}})}$, the method employed above to establish the full row rank of $H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})$} from the full row rank matrix $H_{L}^{cum}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})$} is also applicable to determining the full row rank of $\begin{bmatrix}{H_{L}^{cum}(\{z^{[0,T_{0}-1]}_{i}\}_{i=1}^{\bar{p}})}&{H_{L}^{mos}(\{z^{[0,T_{i}-1]}_{i}\}_{i=\bar{p}+1}^{p})}\end{bmatrix}$} from the full row rank matrix ${H_{L}^{cum}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{\bar{p}}})}+{H_{L}^{cum}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=\bar{p}+1}^{p}})}$. Therefore $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also HCPE.
2). i). If $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are MCPE and $T_{0}=T_{1}=T_{2}=...=T_{p}$, we have ${H_{L}^{cum}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}={H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}({\textbf{1}_{p}}\otimes{I_{{T_{0}}-L+1}})$. Then,
$\text{rank}\ {H_{L}^{cum}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}=\text{rank}\ {H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}$
$~~~-~\text{dim}[\text{im}\,{H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})^{T}}\cap\text{im}({{\bf{1}}_{p}}\otimes{I_{{T_{0}}-L+1}})^{\perp}].$
Thus, we have
$\text{rank}\ {H_{L}^{cum}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}=\text{rank}\ {H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}$
if $\text{im}\,{H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})^{T}}\cap\text{leftker}\,{\textbf{1}_{p}}\otimes{I_{{T_{0}}-L+1}}=\{0\}$, i.e., $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also CCPE.
2). ii). If $T_{0}=T_{1}=T_{2}=...=T_{\bar{p}}$, we have
${H_{L}^{hyb}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}$
$~~~~={H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{\bar{p}}})}(\begin{bmatrix}{{\textbf{1}_{\bar{p}}}}&{{0_{\bar{p}\times(p-\bar{p})}}}\\ {{0_{p-\bar{p}}}}&{{I_{p-\bar{p}}}}\end{bmatrix}\otimes{I_{{T_{0}}-L+1}}).$
Similarly to the above, if additionally, $\text{im}\,{{H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{\bar{p}}})}^{T}}\cap\text{leftker}\,{\begin{bmatrix}{{\textbf{1}_{\bar{p}}}}&{{0_{\bar{p}\times(p-\bar{p})}}}\\ {{0_{p-\bar{p}}}}&{{I_{p-\bar{p}}}}\end{bmatrix}}\otimes{I_{{T_{0}}-L+1}}=\{0\}$, then $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are HCPE.
3). i). If $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are HCPE, then $T_{0}=T_{1}=...=T_{\bar{p}}$ and ${H_{L}^{hyb}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}{H_{L}^{hyb}(\{z^{[0,T_{i}-1]}_{i}\}_{i=1}^{p})}^{T}>0$, i.e., ${M_{1}}+\sum\limits_{i=1}^{\bar{p}}{\left({{\alpha_{i}}{H_{L}}({z_{i}^{[0,{T_{0}}-1]}})\sum\limits_{i\neq j=1}^{\bar{p}}{{\alpha_{j}}H_{L}^{T}({z_{j}^{[0,{T_{0}}-1]}})}}\right)}>0$. Similar to the proof of conclusion 1), we can deduce $M_{1}>0$, and thus $\{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ are also MCPE.
3). ii). If $T_{0}=T_{\bar{p}+1}=...=T_{p}$, then $H_{L}^{hyb}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})$} can be regarded as a special case of $H_{L}^{mos}(\{{z_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})$} in conclusion 2). i). The proof for this case is evidently straightforward. $\blacksquare$
B. Proof of Lemma 1
Proof: We begin by demonstrating that
$\text{rank}\,\begin{bmatrix}{H_{1}^{cum}(\{{x_{i}^{[0,{T_{i}-L}]}\}_{i=1}^{p}})}\\ {H_{L}^{cum}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix}=n+mL,$
i.e., $\begin{bmatrix}{H_{1}^{cum}(\{{x_{i}^{[0,{T_{i}-L}]}\}_{i=1}^{p}})}\\ {H_{L}^{cum}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix}$ achieves full row rank when $\{u_{i}^{[0,T_{i}-1]}\}_{i=1}^{p}$ is CCPE.
[…]
$\begin{bmatrix}{H_{1}^{cum}(\{{x_{i}^{[1,{T_{i}-L-n+1}]}\}_{i=1}^{p}})}\\ {H_{L}^{cum}(\{{u_{i}^{[1,{T_{i}-n}]}\}_{i=1}^{p}})}\end{bmatrix}$=
$\begin{bmatrix}{\begin{bmatrix}A&B\end{bmatrix}}&{{0_{n\times m(n+L-1)}}}\\ {{0_{mL\times(n+m)}}}&{\begin{bmatrix}I_{mL}&{{0}}\end{bmatrix}}\end{bmatrix}\begin{bmatrix}{H_{1}^{cum}(\{{x_{i}^{[0,{T_{i}-L-n}]}\}_{i=1}^{p}})}\\ {H_{L+n}^{cum}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix}$
$\vdots$
$\begin{bmatrix}{H_{1}^{cum}(\{{x_{i}^{[n,{T_{i}-L}]}\}_{i=1}^{p}})}\\ {H_{L}^{cum}(\{{u_{i}^{[n,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix}$=
$\begin{bmatrix}{\begin{bmatrix}{{A^{n}}}&{{A^{n-1}}B}&\cdots&B\end{bmatrix}}&{{0}}\\ {{0_{mL\times n(1+m)}}}&I_{mL}\end{bmatrix}\begin{bmatrix}{H_{1}^{cum}(\{{x_{i}^{[0,{T_{i}-L-n}]}\}_{i=1}^{p}})}\\ {H_{L+n}^{cum}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix}$.
Since $r_{x}\ r_{u}]$ also resides in the left kernel of the matrices on the left side of the above equations, a simultaneous left-multiplication of all equations by $r_{x}\ r_{u}]$ yields
$R_{x}^{u}\begin{bmatrix}{H_{1}^{cum}(\{{x_{i}^{[0,{T_{i}-L-n}]}\}_{i=1}^{p}})}\\ {H_{L+n}^{cum}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix}={0_{(n+1)\times({T_{i}}-L-n+1)}},$
where $R_{x}^{u}=\begin{bmatrix}{{r_{x}}}&{{r_{u}}}&{}&{}&{}\\ {{r_{x}}A}&{{r_{x}}B}&{{r_{u}}}&{}&{}\\ \vdots&\ddots&\ddots&\ddots&{}\\ {{r_{x}}{A^{n}}}&{{r_{x}}{A^{n-1}}B}&\cdots&{{r_{x}}B}&{{r_{u}}}\end{bmatrix}$.
Through CCPE condition, one obtains $\text{rank}\,{H_{L+n}^{cum}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}=m(n+L)$. There must have $\text{dim}(\text{leftker}\begin{bmatrix}{H_{1}^{cum}(\{{x_{i}^{[0,{T_{i}-L-n}]}\}_{i=1}^{p}})}\\ {H_{L+n}^{cum}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix})\leq n$, thus $\text{rank}(R_{x}^{u})\leq n$. Since $R_{x}^{u}$ is a lower triangular block matrix with $n+1$ rows, it follows that at least one of these rows causes $R_{x}^{u}$ to drop rank. Consequently, $r_{u}$ and $r_{x}B$ must both be 0 vectors. At this point, there are two cases for letting $R_{x}^{u}$ unrank to $n$.
1. There exists a nonzero constant $\tau_{0}$ such that $\tau_{0}r_{x}=r_{x}A$. Then we have $r_{x}AB=\tau_{0}r_{x}B=0$, and further recursion yields $r_{x}A^{n-1}B=r_{x}A^{n-2}B\cdots=r_{x}B=0$. The reachability of $(A,B)$ ensures that $r_{x}=0$.
2. letting $r_{x}A^{j}B=0$, j=1,2,..,n1j=1,2,..,n-1 and there exists a nonzero constant $\tau_{j}$ such that $\tau_{j}r_{x}A^{j}=r_{x}A^{j+1}$, in which case the same result is obtained with $r_{x}A^{n-1}B=r_{x}A^{n-2}B\cdots=r_{x}B=0$, and thus $r_{x}=0$.
In summary, $\text{rank}(R_{x}^{u})\leq n$ implies that $r_{x}=r_{u}=0$, and thus the matirx $\begin{bmatrix}{H_{1}^{cum}(\{{x_{i}^{[0,{T_{i}-L}]}\}_{i=1}^{p}})}\\ {H_{L}^{cum}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix}$ has full row rank.
Similarly, employing a comparable approach, we can establish that $\begin{bmatrix}{H_{1}^{mos}(\{{x_{i}^{[0,{T_{i}-L}]}\}_{i=1}^{p}})}\\ {H_{L}^{mos}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix}$ or $\begin{bmatrix}{H_{1}^{hyb}(\{{x_{i}^{[0,{T_{i}-L}]}\}_{i=1}^{p}})}\\ {H_{L}^{hyb}(\{{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}})}\end{bmatrix}$ achieves full row rank when $\{u_{i}^{[0,{T_{i}-1}]}\}_{i=1}^{p}$ is MCPE or HCPE of order $L+n$. We will not reiterate this here. $\blacksquare$
C. Proof of Lemma 3
According to condition ii), there must exists some invertible matrix sequence $M(k)$ such that
$2I-{z^{T}}(k)z(k)=M(k){M^{T}}(k)$
for all $k\geq 0$. Let $A(k)=I-z(k)z^{T}(k)$ and $N(k)=z(k)M(k)$, we have
$A^{T}(k)A(k)-I=-N(k){N^{T}}(k)\leq 0.$
Consider the Lyapunov function $V(k)=x^{T}(k)x(k)$, we can compute the time difference of $V(k)$ as
$V(k+1)-V(k)$ $=-x^{T}(k)(A^{T}(k)A(k)-I)x(k)$
$\leq 0,$
which implies that the equilibrium $x_{e}=0$ is stable in the sense of Lyapunov. Recall the result (Ioannou and Fidan, 2006, Th. A.12.22) of exponential stability, if, further, the pair $(A(k)-K(k)N^{T}(k),N(k))$ is uniformly completely observable (UCO) for any bounded $K(k)$, then $x_{e}=0$ is e.s. Let us choose
$K(k)=-z(k){M^{T}}(k)^{-1},$
which is bounded due to condition i). Then we have $A(k)-K(k)N^{T}(k)=I$. Therefore, it is only necessary to prove the fact that the pair $(I,N(k))$ is UCO.
If $z_{[k,k+l-1]}$ is PE of order $1$ for a fixed constant $l\geq 1$, we have $H_{L}(z_{[k,k+l-1]})H^{T}_{L}(z_{[k,k+l-1]})>0$. Hence, there exists a constant $\beta_{0}>0$ such that
$H_{1}(z_{[k,k+l-1]})H^{T}_{1}(z_{[k,k+l-1]})$ $=\sum\limits_{t=0}^{l-1}{z(k+t)z^{T}(k+t)}$
$\geq{\beta_{0}}I.$ (29)
By (C. Proof of Lemma 3) and condition i), we know that
${\beta_{0}}I\leq\sum\limits_{t=0}^{l-1}{z(k+t){z^{T}}(k+t)}\leq{\beta_{1}}I$ (30)
for some $\beta_{1}\geq\beta_{0}>0$. In addtion, since $N(k){N^{T}}(k)=z(k)(2I-{z^{T}}(k)z(k)){z^{T}}(k)$, one has
$(2-z_{m}^{2})\sum\limits_{t=0}^{l-1}{z(k+t){z^{T}}(k+t)}\leq\sum\limits_{t=0}^{l-1}{N(k+t){N^{T}}(k+t)}$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\leq 2\sum\limits_{t=0}^{l-1}{z(k+t){z^{T}}(k+t)}.$ (31)
Combining (30) and (C. Proof of Lemma 3), we have
${\beta_{0}}(2-z_{m}^{2})I$ $\leq\sum\limits_{t=0}^{l-1}{N(k+t){N^{T}}(k+t)}\leq 2{\beta_{1}}I,$ (32)
which implies that $(I,N(k))$ is UCO and the equilibrium state $x_{e}=0$ is e.s. $\blacksquare$